# Rotation sets and axes in the fine curve graph for torus homeomorphisms

### v1

@@ Abstract.
As a consequence we find new stable criteria for positive scl, and for two homeomorphisms to generate a free group.
@@ 1.1. Rotation sets and algebra
The aim of the present paper concerns the group $\mathrm{Homeo}_{0}(T^{2})$ of homeomorphisms of the two-dimensional torus that are isotopic to the identity. Our results relate the rotation set, a classical invariant from dynamical systems, to the algebra of this group, via its action on the fine curve graph. We will recall the definition of the rotation set in section 2 below; for the moment, let us just say that it is a compact and convex subset of the plane, associated to a given element of our group $\mathrm{Homeo}_{0}(T^{2})$, which is defined up to integer translations.
@@
where the limit always exists by sub-additivity. If the group $G$ is perfect, i.e. generated by the set $\mathcal{C}$ of commutators, then $\ell_{\mathcal{C}}(g)$ and $s\ell_{\mathcal{C}}(g)$ are denoted $\mathrm{cl}(g),\mathrm{scl}(g)$ and called the (stable) commutator length. By Bavard duality, positivity of $\mathrm{scl}(g)$ is equivalent to the existence of a homogeneous quasimorphism $\Phi$ such that $\Phi(g)\neq 0$ ([Bav91]). The identity component of the group of homeomorphisms of any compact manifold is known to be perfect ([And58]), in particular the definition applies to our group $\mathrm{Homeo}_{0}(T^{2})$.
By Kwapisz realization theorem [Kwa92], every polygon with all vertices having rational coordinates (for short, rational polygon) can be realized as the rotation set of some element of $\mathrm{Homeo}_{0}(T^{2})$, and the proof gives an explicit construction as an axiom A diffeomorphism. This provides plenty of examples to illustrate both theorems. Note that a symmetric polygon has an even number of sides; in particular, triangles are never symmetric. The easiest example satisfying Theorem 2 is obtained for a homeomorphism whose rotation set is the triangle $(0,0),(1,0),(0,1)$; we will give a construction close to that of Kwapisz in section 11.2. On the other hand every symmetric quadrilateral $Q$ (parallelogram!) which is rational is the rotation set of some $f$ which has a power conjugate to its inverse, which easily implies $\mathrm{scl}(f)=0$, see section 11.1. This shows some kind of optimality for Theorem 2. By the continuity property of the rotation set, the hypotheses of both theorems are stable: the second theorem describes an open set in $\mathrm{Homeo}_{0}(T^{2})$, and the first one an open set of its cartesian square.
The existence of elements with positive scl is not new: in [BHW22], Bowden, Hensel and Webb proved this in $\mathrm{Homeo}_{0}(\Sigma)$ for every surface $\Sigma$ of genus at least one. This was the original motivation for the introduction of the fine curve graph (see below), the elephant in the room in Theorem 1 and 2. The proof in [BHW22] was rather constructive but did not provide any dynamical feature of the exhibited elements. In [BHW24], the same authors provide more specific examples, namely (relative) pseudo-Anosov maps. We refer the reader to these previous papers for more context and motivation. The following corollary is new.
@@ 1.2. Rotation sets and the fine curve graph
The fine curve graph $\mathcal{C}^{\dagger}(\mathbb{T}^{2})$, introduced by Bowden, Hensel and Webb in [BHW22], is the graph whose vertices are essential simple closed curves, and edges join curves that are either disjoint or have only one intersection point, at which they cross transversely. This graph is Gromov-hyperbolic ([BHW22, Theorem 3.10]). The group $\mathrm{Homeo}_{0}(T^{2})$ acts naturally on this graph, and identifies with its automorphism group ([LMP+21, LRW24]). The translation length $\mathrm{tl}(f)$ of $f\in\mathrm{Homeo}_{0}(T^{2})$ is defined as
@@
We remark that the theorem is still true in higher genus if one replaces the rotation set by the ergodic rotation set ([GM23]), and that both conditions are equivalent to $f$ defining a pseudo-Anosov mapping class relative to a finite set of periodic points. Neither of these are relevant for the current work.
@@
The existence of free groups claimed by Theorem 1 follows easily from theorem 3 by the classical ping-pong argument (see [dlH00]). In section 11 we show how to determine the scaled rotation sets on some interesting examples. In particular, we also get a free group generated by two elements having the same rotation set but distinct scaled rotation sets. By using a different invariant (the twist number, see section 1.4 below), we also get free groups generated by elements sharing the same scaled rotation sets.
To get positive stable commutator length, we need a stronger property. In [BF02], Bestvina and Fujiwara define an equivalence relation on loxodromic elements of the isometry group of a Gromov-hyperbolic space, namely $f\sim g$ whenever $g$ has conjugate elements whose repulsive and attracting fixed points in $\partial_{\infty}X$ are arbitrarily close respectively to the repulsive and attracting fixed points of $f$ (see section 3.4 for more details). The following statement generalizes Theorem 3.
@@
Let $f\in\mathrm{Homeo}_{0}(T^{2})$ whose rotation set has non-empty interior and no point symmetry. Since $\mathrm{Rot}(f^{-1})=-\mathrm{Rot}(f)$, and $\mathrm{tl}(f^{-1})=\mathrm{tl}(f)$, the absence of point symmetry yields that the scaled rotation sets of $f$ and $f^{-1}$ are not translates of each other. From Theorem 4 we get that $f\not\sim f^{-1}$. The result now follows from the work of Bestvina-Fujiwara [BF02], see section 3.4 below. ∎
@@ 1.3. The metric WPD property
The heart of our argument is the discovery that loxodromic elements of the fine curve graph share some kind of weak proper discontinuity property that we describe now. This property is a natural extension of Bestvina-Fujiwara WPD property ([BF02], see also [Osi16]) to the setting of Rosendal’s geometric group theory [Ros21], which enlarges geometric group theory to enclose some non locally compact groups.
@@
Mann-Rosendal’s theorem states that for every compact manifold, the identity component of the group of homeomorphisms is locally (OB), and the fragmentation norm with respect to any given cover by open balls is quasi-isometric to the maximal metric [MR18]. Back to our context, we propose to use the following concrete incarnation of the maximal metric. For $f,g\in\mathrm{Homeo}_{0}(T^{2})$, define
@@
Our theorems will all be consequences of the following proposition.
Proposition 1.3.
@@
Let $f,g\in\mathrm{Homeo}_{0}(T^{2})$ be acting loxodromically on the fine curve graph. Then $\xi^{+}(f)=\xi^{+}(g)$ if and only if there exists a sequence of positive numbers $(m_{n})_{n\geq 0}$ such that the sequence $(f^{-n}g^{m_{n}})_{n\geq 0}$ is bounded in the group $\mathrm{Homeo}_{0}(T^{2})$.
Theorem 3 above is actually an easy consequence of this corollary. Let us describe another dynamical consequence. Given $f\in\mathrm{Homeo}_{0}(T^{2})$, consider the equivalence relation on the torus defined by
$x\sim_{f^{-}}y\Leftrightarrow\text{ the sequence }(\|\widetilde{f}^{n}(\widetilde{x})-\widetilde{f}^{n}(\widetilde{y})\|)_{n\leq 0}\text{ is bounded}$
where the norm is the sup norm in $\mathbb{R}^{2}$, $\widetilde{f}$ denotes any lift of $f$, and $\widetilde{x},\widetilde{y}$ any lifts of $x,y$ (the property does not depend on the choices of the lifts). We call the partition by equivalence classes the backward shadowing partition. Note that all the points of a given class share the same backward rotation property, e.g. if any point admits a rotation vector then all the other points admit the same rotation vector (again, see section 2 for the definitions). The following corollary is immediate.
@@
Let $f,g\in\mathrm{Homeo}_{0}(T^{2})$ be acting loxodromically on the fine curve graph, and assume $\xi^{+}(f)=\xi^{+}(g)$. Then the backward shadowing partitions of $f$ and $g$ coincide. Any $h\in\mathrm{Homeo}_{0}(T^{2})$ that fixes $\xi^{+}(f)$ preserves the relation $\sim_{f^{-}}$, that is, every image of an equivalence class is an equivalence class.
In the previous corollary, it may seems paradoxical that the attracting fixed point on the fine curve graph is related to the backward shadowing relation on the torus. This comes from the fact that the action on the fine curve graph is a left action, whereas the natural metric on the group is right invariant. Namely, if we have a bound, say, on the $C^{0}$ distance between $f^{-n}$ and $g^{-m}$ for large positive numbers $m$ and $n$, then we have the same bound on $f^{-n}g^{m}$, this will translate into a bound in the fine curve graph between a given curve $\alpha$ and $f^{-n}g^{m}\alpha$, and thus also between $f^{n}\alpha$ and $g^{m}\alpha$.
Inhyeok Choi has independently proved that relative pseudo-Anosov homeomorphisms have metric WPD in every genus, see [Cho25]. His proof uses very different methods than our approach, using an explicit description of the fixed points at infinity as singular foliations on the surface from [BHW24].
@@ 1.4. Ingredients for metric WPD
Inspired by the marking graph hierarchy formula of Masur-Minsky [MM00] in the classical curve graph setting, we get an upper bound for the distance in the fine marking graph, in terms of (1) the distance in the fine curve graph between the curves corresponding to the sides of the fine markings, and (2) the sum of the twist numbers in suitable annuli (Lemma 8.1). Here the twist number in an annulus $A$ between two curves that cross $A$ is the distance in the fine arc graph of $A$ between the pieces of the curves that cross $A$. In other words, this is the fine version, in genus 1, of subsurface projections. Another key ingredient is a fine version of the Bounded Geodesic Image Theorem (Lemmas 6.4 and 6.5), from which we deduce that twist numbers between two curves along a given loxodromic axis are bounded independently of the annulus where we project (Lemma 6.7). The WPD property is then an easy consequence of these ingredients.
@@
A first proof of Theorem 3 involved train tracks via Masur-Minsky nesting lemma and the fact that iterates of a given curve determines the rotation set (see [Le 24]). We were not able to prove positive stable commutator length along the same lines.
@@ 1.5. Organization of the paper
Sections 23 and 4 are preliminaries dealing respectively with the rotation set, hyperbolic geometry and the fine curve graph. The fine marking graph and its relation with the group $\mathrm{Homeo}(T^{2})$ is described in section 5. In section 6 we introduce the fine twist numbers, prove the bounded geodesic image theorems, and deduce that twist numbers are bounded along loxodromic axes. In Section 7 we discuss a variant of the fine curve graph and prove technical distance bounds which are the main ingredient for a distance formula in the fine marking graph that is proved in Section 8. We prove the metric WPD property in section 9. The main theorems are proved in section 10. The last section 11 is dedicated to describing several examples, and also contains the definition of the twist number of a homeomorphism.
@@ 2.1. Definition
We recall the definition and main properties of the rotation set, following [MZ90]. Here the torus is $T^{2}=\mathbb{R}^{2}/\mathbb{Z}^{2}$. Let $f\in\mathrm{Homeo}_{0}(T^{2})$, and fix a lift $\widetilde{f}$ of $f$: this is an element of the group $\mathrm{Homeo}(\mathbb{R}^{2},\mathbb{Z}^{2})$ of homeomorphisms of the plane that commutes with integer translations. Consider the function D(f~)D($\widetilde${f}}):
@@
Theorem (Misiurewicz-Ziemian, [MZ90]).
@@ 3. Hyperbolic preliminaries
In this section, we collect some preliminaries on Gromov hyperbolic spaces. We refer the reader to e.g. [BH99] for details on hyperbolic spaces; we will only collect the required results. Most of them are well-known, but especially in Subsections 3.2 or 3.3 we discuss results specifically adapted to our situation.
@@ 3.1. Basic definitions and the boundary
If $X$ is proper, then $\partial_{\infty}X$ can be identified with the set of geodesic rays up to bounded distance – but since the spaces we deal with are very much not proper, it is not clear if every boundary point is reached by an actual geodesic ray in our setting. See e.g. [LT25, Section 2.A] for discussion.
@@ 3.2. Geodesics and quasigeodesics
A central property of geodesics in hyperbolic spaces is the following stability (see e.g. [BH99, Chapter III.H, Theorem 1.7])

### v2

@@ Abstract.
As a consequence we find new stable criteria for positive scl, and for two homeomorphisms to generate a free group, and we provide a Tits alternative for groups of torus homeomorphisms.
@@ 1.1. Rotation sets and algebra
The aim of the present paper concerns the group $\mathrm{Homeo}_{0}(T^{2})$ of homeomorphisms of the two-dimensional torus that are isotopic to the identity. Our results relate the rotation set, a classical invariant from dynamical systems, to the algebra of this group, via its action on the fine curve graph. We will recall the definition of the rotation set in section 2 below; for the moment, let us just say that it is a compact and convex subset of the plane, associated to a given element of our group $\mathrm{Homeo}_{0}(T^{2})$, which is defined up to integer translations. We state three results relating the rotation set to the algebra of the group. The first theorem is a kind of Tits alternative; the second gives a criterion for two elements to generate a free group; the third one for positive stable commutator length.
Theorem 1.
Let $G$ be a subgroup of $\mathrm{Homeo}_{0}(T^{2})$ which contains some element $f$ whose rotation set $\mathrm{Rot}(f)$ has interior. Then one of the following possibilities occurs:
  1. (1)
    $G$ contains a free group, generated by some power $f^{n}$ of $f$ and some conjugate of $f^{n}$;
  2. (2)
    $G$ or a subgroup of index $2$ of $G$ fixes every backward shadowing equivalence class of $f$, or every forward shadowing class of $f$.
The forward shadowing class of a point $x$ is the set of points whose positive orbit stays a bounded distance from the positive orbit of $f$ in the universal cover. This is a non-trivial equivalence relation, in the sense that it contains uncountably many distinct equivalence classes, see section 11 below for more details.
We mention that there are other Tits alternatives for homeomorphism groups of surfaces [HX, Dagger3], which assume much stronger conditions on an element of the group – namely being a (pseudo-)Anosov – but also find stronger constraints in the case where no free group is found – namely fixing a (singular) foliation up to index $2$.
@@
where the limit always exists by sub-additivity. If the group $G$ is perfect, i.e. generated by the set $\mathcal{C}$ of commutators, then $\ell_{\mathcal{C}}(g)$ and $s\ell_{\mathcal{C}}(g)$ are denoted $\mathrm{cl}(g),\mathrm{scl}(g)$ and called the (stable) commutator length. By Bavard duality, positivity of $\mathrm{scl}(g)$ is equivalent to the existence of a homogeneous quasimorphism $\Phi$ such that $\Phi(g)\neq 0$ ([bavard]). The identity component of the group of homeomorphisms of any compact manifold is known to be perfect ([Anderson]), in particular the definition applies to our group $\mathrm{Homeo}_{0}(T^{2})$.
By Kwapisz realization theorem [kwapisz_realise], every polygon with all vertices having rational coordinates (for short, rational polygon) can be realized as the rotation set of some element of $\mathrm{Homeo}_{0}(T^{2})$, and the proof gives an explicit construction as an axiom A diffeomorphism. This provides plenty of examples to illustrate both theorems. Note that a symmetric polygon has an even number of sides; in particular, triangles are never symmetric. The easiest example satisfying Theorem 3 is obtained for a homeomorphism whose rotation set is the triangle $(0,0),(1,0),(0,1)$; we will give a construction close to that of Kwapisz in section 12.2. On the other hand every symmetric quadrilateral $Q$ (parallelogram!) which is rational is the rotation set of some $f$ which has a power conjugate to its inverse, which easily implies $\mathrm{scl}(f)=0$, see section 12.1. This shows some kind of optimality for Theorem 3. Also note that by the continuity property of the rotation set, Theorem 3 describes an open set in $\mathrm{Homeo}_{0}(T^{2})$, and Theorem 2 an open set of its cartesian square.
The existence of elements with positive scl is not new: in [Dagger1], Bowden, Hensel and Webb proved this in $\mathrm{Homeo}_{0}(\Sigma)$ for every surface $\Sigma$ of genus at least one. This was the original motivation for the introduction of the fine curve graph (see below), the elephant in the room in Theorem 2 and 3. The proof in [Dagger1] was rather constructive but did not provide any dynamical feature of the exhibited elements. In [Dagger3], the same authors provide more specific examples, namely (relative) pseudo-Anosov maps. We refer the reader to these previous papers for more context and motivation. The following corollary is new.
@@ 1.2. Rotation sets and the fine curve graph
The fine curve graph $\mathcal{C}^{\dagger}(\mathbb{T}^{2})$, introduced by Bowden, Hensel and Webb in [Dagger1], is the graph whose vertices are essential simple closed curves, and edges join curves that are either disjoint or have only one intersection point, at which they cross transversely. This graph is Gromov-hyperbolic ([Dagger1, Theorem 3.10]). The group $\mathrm{Homeo}_{0}(T^{2})$ acts naturally on this graph, and identifies with its automorphism group ([LMPVY, LW]). The translation length $\mathrm{tl}(f)$ of $f\in\mathrm{Homeo}_{0}(T^{2})$ is defined as
@@
We remark that the theorem is still true in higher genus if one replaces the rotation set by the ergodic rotation set ([GM]), and that both conditions are equivalent to $f$ defining a pseudo-Anosov mapping class relative to a finite set of periodic points. Neither of these are relevant for the current work.
@@
The existence of free groups claimed by Theorem 2 follows easily from theorem 4 by the classical ping-pong argument (see [delaHarpe]). In section 12 we show how to determine the scaled rotation sets on some interesting examples. In particular, we also get a free group generated by two elements having the same rotation set but distinct scaled rotation sets. By using a different invariant (the twist number, see section 1.4 below), we also get free groups generated by elements sharing the same scaled rotation sets.
To get positive stable commutator length, we need a stronger property. In [BF], Bestvina and Fujiwara define an equivalence relation on loxodromic elements of the isometry group of a Gromov-hyperbolic space, namely $f\sim g$ whenever $g$ has conjugate elements whose repulsive and attracting fixed points in $\partial_{\infty}X$ are arbitrarily close respectively to the repulsive and attracting fixed points of $f$ (see section 3.4 for more details). The following statement generalizes Theorem 4.
@@
Let $f\in\mathrm{Homeo}_{0}(T^{2})$ whose rotation set has non-empty interior and no point symmetry. Since $\mathrm{Rot}(f^{-1})=-\mathrm{Rot}(f)$, and $\mathrm{tl}(f^{-1})=\mathrm{tl}(f)$, the absence of point symmetry yields that the scaled rotation sets of $f$ and $f^{-1}$ are not translates of each other. From Theorem 5 we get that $f\not\sim f^{-1}$. The result now follows from the work of Bestvina-Fujiwara [BF], see section 3.4 below. ∎
@@ 1.3. The metric WPD property
The heart of our argument is the discovery that loxodromic elements of the fine curve graph share some kind of weak proper discontinuity property that we describe now. This property is a natural extension of Bestvina-Fujiwara WPD property ([BF], see also [osin]) to the setting of Rosendal’s geometric group theory [Rosendal], which enlarges geometric group theory to enclose some non locally compact groups.
@@
Mann-Rosendal’s theorem states that for every compact manifold, the identity component of the group of homeomorphisms is locally (OB), and the fragmentation norm with respect to any given cover by open balls is quasi-isometric to the maximal metric [MR]. Back to our context, we propose to use the following concrete incarnation of the maximal metric. For $f,g\in\mathrm{Homeo}_{0}(T^{2})$, define
@@
Our theorems will all be consequences of the following one.
Theorem 6.
@@
Let $f,g\in\mathrm{Homeo}_{0}(T^{2})$ be acting loxodromically on the fine curve graph. Then $\xi^{-}(f)=\xi^{-}(g)$ if and only if there exists a sequence of positive numbers $(m_{n})_{n\geq 0}$ such that the sequence $(f^{n}g^{-m_{n}})_{n\geq 0}$ is bounded in the group $\mathrm{Homeo}_{0}(T^{2})$.
Theorem 4 above is actually an easy consequence of this corollary.
The following corollaries provides dynamical interpretation; the last one is the key to our Tits alternative.
@@
Let $f,g\in\mathrm{Homeo}_{0}(T^{2})$ be acting loxodromically on the fine curve graph. Assume $\xi^{-}(f)=\xi^{-}(g)$. Then the forward shadowing classes of $f$ and $g$ coincide.
Corollary 1.5.
Let $f\in\mathrm{Homeo}_{0}(T^{2})$ be acting loxodromically on the fine curve graph. Let $h\in\mathrm{Homeo}_{0}(T^{2})$, and choose lifts $\widetilde{f},\widetilde{h}$ of $h$ and $f$ to the universal cover. Then the following are equivalent:
  1. (1)
    $h(\xi^{-}(f))=\xi^{-}(f)$,
  2. (2)
    the sequence of sup norms
    $\left(\|\widetilde{f}^{n}-\widetilde{f}^{n}\widetilde{h}\|_{\infty}\right)_{n\geq 0}$
    is bounded.
Under these conditions, $h$ preserves each forward shadowing class of $f$.
In the previous corollaries, it may seems paradoxical that the repulsive fixed point on the fine curve graph is related to the forward shadowing relation on the torus. This comes from the fact that the action on the fine curve graph is a left action, whereas the natural metric on the group is right invariant. Namely, if we have a bound, say, on the $C^{0}$ distance between $f^{n}$ and $g^{m}$ for large positive numbers $m$ and $n$, then we have the same bound on $f^{n}g^{-m}$, this will translate into a bound in the fine curve graph between a given curve $\alpha$ and $f^{n}g^{-m}\alpha$, and thus also between $f^{-n}\alpha$ and $g^{-m}\alpha$.
Inhyeok Choi has independently proved that relative pseudo-Anosov homeomorphisms have metric WPD in every genus, see [Choi]. His proof uses very different methods than our approach, using an explicit description of the fixed points at infinity as singular foliations on the surface from [Dagger3].
@@ 1.4. Ingredients for metric WPD
Inspired by the marking graph hierarchy formula of Masur-Minsky [MM2] in the classical curve graph setting, we get an upper bound for the distance in the fine marking graph, in terms of (1) the distance in the fine curve graph between the curves corresponding to the sides of the fine markings, and (2) the sum of the twist numbers in suitable annuli (Lemma 8.1). Here the twist number in an annulus $A$ between two curves that cross $A$ is the distance in the fine arc graph of $A$ between the pieces of the curves that cross $A$. In other words, this is the fine version, in genus 1, of subsurface projections. Another key ingredient is a fine version of the Bounded Geodesic Image Theorem (Lemmas 6.4 and 6.5), from which we deduce that twist numbers between two curves along a given loxodromic axis are bounded independently of the annulus where we project (Lemma 6.7). The WPD property is then an easy consequence of these ingredients.
@@
A first proof of Theorem 4 involved train tracks via Masur-Minsky nesting lemma and the fact that iterates of a given curve determines the rotation set (see [Videos]). We were not able to prove positive stable commutator length along the same lines.
@@ 1.5. Organization of the paper
Sections 23 and 4 are preliminaries dealing respectively with the rotation set, hyperbolic geometry and the fine curve graph. The fine marking graph and its relation with the group $\mathrm{Homeo}(T^{2})$ is described in section 5. In section 6 we introduce the fine twist numbers, prove the bounded geodesic image theorems, and deduce that twist numbers are bounded along loxodromic axes. In Section 7 we discuss a variant of the fine curve graph and prove technical distance bounds which are the main ingredient for a distance formula in the fine marking graph that is proved in Section 8. We prove the metric WPD property in section 9. The theorems relating the rotation set and the axis are proved in section 10. The forward shadowing relation is studied in section 11, which also contains the proof of the Tits alternative. The last section 12 is dedicated to describing several examples, and also contains the definition of the twist number of a homeomorphism.
@@ 2.1. Definition
We recall the definition and main properties of the rotation set, following [MZ]. Here the torus is $T^{2}=\mathbb{R}^{2}/\mathbb{Z}^{2}$. Let $f\in\mathrm{Homeo}_{0}(T^{2})$, and fix a lift $\widetilde{f}$ of $f$: this is an element of the group $\mathrm{Homeo}(\mathbb{R}^{2},\mathbb{Z}^{2})$ of homeomorphisms of the plane that commutes with integer translations. Consider the function D(f~)D($\widetilde${f}}):
@@
Theorem (Misiurewicz-Ziemian, [MZ]).
@@ 3. Hyperbolic preliminaries
In this section, we collect some preliminaries on Gromov hyperbolic spaces. We refer the reader to e.g. [BH] for details on hyperbolic spaces; we will only collect the required results. Most of them are well-known, but especially in Subsections 3.2 or 3.3 we discuss results specifically adapted to our situation.
@@ 3.1. Basic definitions and the boundary
If $X$ is proper, then $\partial_{\infty}X$ can be identified with the set of geodesic rays up to bounded distance – but since the spaces we deal with are very much not proper, it is not clear if every boundary point is reached by an actual geodesic ray in our setting. See e.g. [LongTan, Section 2.A] for discussion.
@@ 3.2. Geodesics and quasigeodesics
A central property of geodesics in hyperbolic spaces is the following stability (see e.g. [BH, Chapter III.H, Theorem 1.7])